A jar contains 19 red marbles, of which 5 are large and 14 are small, and 19 blue marbles, of which 7 are large and 12 are small. If 1 marble is randomly selected from the jar,

a) What is the probability of obtaining a red marble? (Use 2 decimal places.)

(b)What is the probability of obtaining a small marble? (Use 2 decimal places.)

c)What is the probability of obtaining a small red marble? (Use 2 decimal places.)
Answer #1

Ans:

Total marbles=19+19=38

a)

P(red marble)=19/38=0.50

b)

P(small marble)=(14+12)/38=26/38=0.68

c)

P(small red marble)=14/38=0.37